# PillarTrack: Redesigning Pillar-based Transformer Network for Single Object Tracking on Point Clouds

### v1

@@ V Experiments
Figure 6: Visualization results comparison of our PillarTrack and baseline method (SMAT[16]) on the categories of Car, Pedestrian and Cyclist in KITTI dataset.
To comprehensively evaluate the performance of our method, particularly its capability to handle point clouds of varying sparsity, we conduct visualization on the different category, encompassing both sparse and dense scenes. Through comparative experiments on the initial frame and subsequent sequences, our method demonstrated superior tracking accuracy in both sparse and dense point cloud environments compared to the baseline SMAT method, highlighting ours enhanced adaptability and robustness. In the Cyclist category, our method was able to consistently track the target, whereas SMAT occasionally lost the target. For the Pedestrian category, our method maintained stable tracking from the outset, while SMAT failed to accurately identify the target initially. The visualization results in Figure 5 further substantiate the effectiveness and superior performance of our method across various scenarios.
V-D Running Speed
We calculated the average running time of all test frames for car to measure PillarTrack’s speed. PillarTrack model achieved 27 FPS on a single NVIDIA 3080 GPU, while baseline model SMAT has a 18.2 FPS in the inference phase. We did not introduce extra computational parameters, and we further reduced the computational overhead on the backbone while increasing the model’s performance. This once again demonstrates the effectiveness of our PillarTrack. Furthermore, our proposed PillarTrack can leverage quantization techniques[61, 62] to accelerate the runtime. We leave this problem for future research.
@@ VI CONCLUSIONS
In this work, we proposed PillarTrack, a pillar-based 3D single-object tracking framework that decreases computation overhead while improving performance. The framework introduces a Pyramid-type Encoding Pillar Feature Encoder (PE-PFE) and a modality-aware Transformer-based backbone, and a simply replacement of activation function. Our extensive experiments on the KITTI dataset demonstrate PillarTrack’s superior performance, achieving an better trade-off between speed and accuracy. Based on above designs and its efficiency, We hope to inspire the community to rethink the design of network for 3D single object tracking based on point clouds.

### v2

@@ V-B Ablation Study
To comprehensively evaluate the performance of our method, particularly its more representative capability of point clouds, we conduct visualization on the different category, encompassing both sparse and dense scenes. Through comparative experiments on the initial frame and subsequent sequences, our method demonstrated superior tracking accuracy in both sparse and dense scenes compared to the baseline[16] method, highlighting ours enhanced adaptability and robustness. In the Cyclist category, our method was able to consistently track the target, whereas SMAT lost the target. For the Pedestrian category, our method maintained stable tracking from the outset, while SMAT failed to identify the targets. The visualization results in Figure 6 substantiate the effectiveness and superior performance of our method across various scenarios.
V-D Running Speed.
We calculated the average running time of all test frames for car to measure PillarTrack’s speed. PillarTrack model achieved 27 FPS on a single NVIDIA 3080 GPU, while baseline model SMAT only has 18.2 FPS. We did not introduce extra computational parameters, and we further reduced the computational overhead on the backbone while increasing the model’s performance. This once again demonstrates the effectiveness of our PillarTrack. Furthermore, our PillarTrack can leverage quantization techniques[44] to accelerate the runtime. We leave this problem for the future research.
@@ VI CONCLUSIONS
In this work, we proposed PillarTrack, a pillar-based 3D single-object tracking framework that decreases computation overhead while improving performance. We propose a Pyramid-Encoding Pillar Feature Encoder (PE-PFE) and a modality-aware Transformer-based backbone to enhance the pillar representation. Extensive experiments on the KITTI and nuScenes demonstrate the superior performance of PillarTrack. Based on above designs and its efficiency, We hope our method could provide new insights for the design of 3D SOT network based on point clouds.